OpenStudy (anonymous):

Is the following statement true or false (a^4)^(1/4)=a for every real number a.

OpenStudy (anonymous):

\[\sqrt[4]{a^4}=a\]

OpenStudy (anonymous):

What happens if you put in a negative number?

OpenStudy (anonymous):

you get a complex solution.

OpenStudy (anonymous):

or you get a?... idk

OpenStudy (anonymous):

Right about the complex solution, so that means it's not valid for EVERY real number a.

OpenStudy (anonymous):

Err, wait, sorry

OpenStudy (anonymous):

Yeah I tried it and got a... or the number...

OpenStudy (anonymous):

Take a=-1 \[^4\sqrt{(-1)^4}=^4\sqrt{1}=1\]

OpenStudy (anonymous):

But 1 does not equal -1.

OpenStudy (anonymous):

The fourth root of 1 is 1, because 1^4=1

OpenStudy (anonymous):

if 1 is not equal to -1... then a is not equal to a and there is no solution?

OpenStudy (anonymous):

Close. \[a \neq ^4\sqrt{a^4}\] for negative numbers

OpenStudy (anonymous):

Okay, Thank you!!!